OpenStudy (anonymous):

how do I prove sin^2x/-cot^2x

OpenStudy (anonymous):

can you repost the question, it seems bit incomplete

hero (hero):

Not enough information to prove.

OpenStudy (anonymous):

I'm sorry it is supposed to be set equal to -sin^2x - tan^2x

hero (hero):

\[-\frac{\sin^2x}{\cot^2x} = -\sin^2x - \tan^2x\] False